OpenStudy (anonymous):

PLEASE HELP!!!!!! Could you please find the answer and explain how you got it? Thanks!: If f(x) = 6^x and g(x) = log6 x, find f(g(x)).

jimthompson5910 (jim_thompson5910):

f(x) = 6^x f(g(x)) = 6^(g(x)) f(g(x)) = 6^( log6 x ) f(g(x)) = x Note: Keep in mind that b^(log base b of x) = x

OpenStudy (anonymous):

Thanks so much, jim_thompson5910! I appreciate it!

jimthompson5910 (jim_thompson5910):

you're welcome